# On the comparison of ultradifferentiable weight function settings

### v1

@@ 5.1. On the exponential-type case
    (4.9) holds if and only if $\sup_{x\in\mathbb{R}}\omega(x)=+\infty$ and (4.10) if and only if $\lim_{|x|\rightarrow+\infty}\omega(x)=+\infty$.
@@ 5.2. On the dilatation-type case
    (4.9) holds if and only if $\sup_{x\in\mathbb{R}}\omega(ax)-\omega(x)=+\infty$ for some $a>1$ and (4.10) if and only if $\lim_{|x|\rightarrow+\infty}\omega(ax)-\omega(x)=+\infty$ for some $a>1$.
@@
    If $\omega$ is non-decreasing and radial, then (4.14) resp. (4.15) for $\mathcal{W}^{\mathfrak{c}}_{\omega}$ amounts to have $\omega(\ell_{1}(n+1))\leq\omega(\ell n)+C$ for some $C\geq 1$ and all $n\in\mathbb{N}$. Thus, both properties hold automatically when $\omega$ is assumed to be non-decreasing: In the first estimate choose $\ell_{1}:=\frac{\ell}{2}$ and in the second one $\ell:=2\ell_{1}$.
@@ 6.1. Weighted subspaces of a fixed Banach space
    Section 3.4: By the analogous definition and system of seminorms and by recalling that the proof of Proposition 3.6 purely involves weights and their growth properties, it follows that this result applies to $\mathcal{A}^{p}$.
@@
    Section 4.1; main Theorems 4.1 and 4.2: Since the spaces are expressed by the analogous system of seminorms and by taking into account the previous comment $(c)$ we have that the crucial results Propositions 4.3, 4.5 and 4.6 hold in this setting, too. Note (again) that the proofs of Propositions 4.5 and 4.6 require the non-triviality of the corresponding (smaller) weighted classes and which is not clear by comment $(b)$ above. However, when taking into account the conventions mentioned in Remark 4.4 one is able to conclude in any case.
@@ 6.2. Weighted Fourier image spaces
    Assume that $\{0\}\neq\widehat{\mathcal{A}}^{p}_{[\mathcal{S}]}$ holds, then the proofs of Propositions 4.5 and 4.6 can be transferred as follows:
    Take a (fixed) non-trivial $f\in\widehat{\mathcal{A}}^{p}_{[\mathcal{S}]}$ and for any $x_{0}\in\mathbb{R}^{d}$ put $g_{x_{0}}(x):=e^{i\langle x,x_{0}\rangle}f(x)$. Then $\widehat{g}_{x_{0}}(x)=\widehat{f}(x-x_{0})$ and apply (4.3), (4.4), (4.5), and (4.6) to the family $\{g_{x_{0}}:x_{0}\in\mathbb{R}^{d}\}$; see also the original proof of [3, Thm. 1.3.18] (for the Beurling-type).
@@
    Indeed, Proposition 3.6 transfers and shows that under (3.13) resp. (3.14) the corresponding weighted classes are modulation invariant.
@@ 6.3. Testfunction spaces appearing in the literature
    Finally, in [23] for $\omega$ satisfying conditions (1)(4)(1)-(4) from Section 2.8 the following test function classes have been studied for $U\subseteq\mathbb{R}^{d}$ (see [23, p. 18-19]):
@@
Now focus on the exponential-type system and first note that the equivalences stated in $(c)$ and $(d)$ in Section 5.1 are not valid anymore in this situation: Recall that $(\omega_{1})$ for a radial $\omega$ means $\omega(2t)\leq L\omega(t)+L$ for some $L\geq 1$ and all $t\geq 0$. This constant $L$ yields then a uniform scaling in the parameters of the matrix $\mathcal{W}^{\mathfrak{c}}_{\omega}$; i.e. one infers the relation $L\ell=n$ and which is a problem when restricting the parameters and when $\ell\rightarrow\Lambda^{\prime}$ (resp. $n\rightarrow\Lambda$). Indeed, $L$ yields a transformation (dilation) of the truncation parameter.
@@ 7. On the failure of the convexity condition $(\omega_{4})$
    By (I)(III)(I)-(III), and in view of Remark 2.5 and Lemmas 2.3 and 2.6, we have that $\omega$ is equivalent to a weight function in the sense of Beurling-Björck (see Definition 2.10).
@@
When combining the information from $(\mathfrak{C}_{1})$ and $(\mathfrak{C}_{2})$ with the characterizations obtained in Theorems 4.1 and 4.2 (see also Proposition 4.9), and by taking into account the comments in Section 5.1, especially the characterization of the growth relations in comments $(c)$ and $(d)$ there, within the exponential-type weight function matrix setting we obtain: There exist examples of weighted classes treated in the Beurling-Björck-setting which cannot be described (alternatively) by any BMT-weight function and this statement applies to both Roumieu- and Beurling-type spaces.
@@ 7.2. Construction of the (counter-)example
Let $\delta=(\delta_{j})_{j\in\mathbb{N}_{>0}}$ be an admissible parameter sequence, arbitrary but from now on fixed. The definition of $\omega$ is depending on chosen $\delta$ which suggests the notation $\varphi_{\omega_{\delta}}$. However, since $\delta$ is now fixed, in order to lighten notation we skip $\delta$ and simply write $\varphi_{\omega}$ except in the very last part Step g: There we verify $(V)$ and hence we have to emphasize the different parameter sequences.
@@
(7.11) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;\varphi_{\omega}(y_{j}):=\frac{2j\varphi_{\omega}(x_{j})}{1-jt_{j}}=\frac{2\delta_{j}j^{2}x_{j}}{1-jt_{j}},$
@@
and let us see that this estimate holds: First, $x_{j}\frac{1-jt_{j}}{2}<\frac{x_{j}}{2}$ is clear and so it suffices to have $\left(j-\frac{1}{2}\right)x_{j}\geq j\frac{1-jt_{j}}{2t_{j}}$. Second, by taking into account (7.15) for this estimate it suffices to have $\left(j-\frac{1}{2}\right)\frac{2j}{t_{j}}\geq\frac{j}{2t_{j}}$ which is equivalent to having $4j^{2}-2j\geq j$, i.e. to $\frac{4}{3}j\geq 1$ which is clear for all $j\in\mathbb{N}_{>0}$. Thus the claim is verified.
@@
Similarly, for $\ell_{j}$ one has by definition that
$\ell_{j}$ $:=\frac{\varphi_{\omega}(x_{j+1})-\varphi_{\omega}(y_{j})}{x_{j+1}-y_{j}}=\frac{2\delta_{j+1}j(j+1)^{2}x_{j}(1-jt_{j})^{-1}-2\delta_{j}j^{2}x_{j}(1-jt_{j})^{-1}}{2j(j+1)x_{j}(1-jt_{j})^{-1}-2jx_{j}(1-jt_{j})^{-1}}$
$=\frac{\delta_{j+1}(j+1)^{2}-\delta_{j}j}{(j+1)-1}=\frac{\delta_{j+1}(j+1)^{2}-\delta_{j}j}{j}.$
@@
and (7.10) and (7.11) together give
$\frac{\varphi_{\omega}(y_{j})}{y_{j}}=\frac{2\delta_{j}j^{2}x_{j}}{1-jt_{j}}\frac{1-jt_{j}}{2jx_{j}}=\delta_{j}j,$
@@
then the corresponding weights $\omega_{\delta}$ and $\omega_{\delta^{\prime}}$ are not equivalent.
@@
Let an admissible parameter sequence $\delta$ satisfying $\lim_{j\rightarrow+\infty}\delta_{j}=0$ be given; consider e.g. $\delta_{j}:=\frac{1}{\log(e+j)}$. However, note that this additional condition is not required necessarily in the above arguments and $\delta$ can be bounded by below by some $\epsilon>0$. In any case, under this assumption each power sequence $\delta^{\alpha}=(\delta_{j}^{\alpha})_{j\in\mathbb{N}_{>0}}$, $0<\alpha<1$, is again an admissible parameter sequence since $\delta_{j+1}^{\alpha}\leq\delta_{j}^{\alpha}\leq(j+2)^{\alpha}\delta_{j+1}^{\alpha}\leq(j+2)\delta_{j+1}^{\alpha}$ for all $j\in\mathbb{N}_{>0}$ and $\delta_{j}^{\alpha}\geq\delta_{j}$ for all $j$ such that $\delta_{j}\leq 1$; i.e. for all $j$ sufficiently large. Thus all requirements from (7.5) hold for $\delta^{\alpha}:=(\delta^{\alpha}_{j})_{j\in\mathbb{N}_{>0}}$ and (each) $\delta^{\prime}:=\delta^{\alpha}$ serves as an example for an admissible parameter sequence satisfying the second condition in (7.21).

### v2

@@ 5.1. On the exponential-type case
    (4.11) holds if and only if $\sup_{x\in\mathbb{R}}\omega(x)=+\infty$ and (4.12) if and only if $\liminf_{|x|\rightarrow+\infty}\omega(x)=+\infty$.
@@ 5.2. On the dilatation-type case
    (4.11) holds if and only if $\sup_{x\in\mathbb{R}}\omega(ax)-\omega(x)=+\infty$ for some $a>1$ and (4.12) if and only if $\liminf_{|x|\rightarrow+\infty}\omega(ax)-\omega(x)=+\infty$ for some $a>1$.
@@
    If $\omega$ is radial, then (4.16) resp. (4.17) for $\mathcal{W}_{\omega,\mathfrak{c}}$ amounts to have $\omega(\ell_{1}(n+1))\leq\omega(\ell n)+C$ for some $C\geq 1$ and all $n\in\mathbb{N}$. Thus, both properties hold automatically when $\omega$ is also assumed to be non-decreasing: In the first estimate choose $\ell_{1}:=\frac{\ell}{2}$ and in the second one $\ell:=2\ell_{1}$.
@@ 6.1. Weighted subspaces of a fixed Banach space
    Section 3.4: The proof of Proposition 3.6 purely involves weights and their growth properties and transfers immediately. However, when $\mathcal{A}^{p}$ is not translation invariant, then it is not clear in general that for given $f\in\mathcal{A}^{p}$ we get $T_{x_{0}}(f)\in\mathcal{A}^{p}$ and hence by definition the estimates in Proposition 3.6 do not yield the fact that (each) $T_{x_{0}}$ is acting continuously between weighted spaces $\mathcal{A}^{p}_{[\mathcal{T}]}$ and $\mathcal{A}^{p}_{[\mathcal{S}]}$. Therefore, it is natural to assume from the beginning that $\mathcal{A}^{p}$ is translation invariant and then this result immediately transfers to the weighted $\mathcal{A}^{p}$-setting.
@@
    Section 4.1; main Theorems 4.1 and 4.2: Proposition 4.3 is clear by the analogous system of seminorms. And the proofs of the crucial Propositions 4.5 and 4.6 transfer when $\mathcal{A}^{p}$ is translation invariant by taking into account the previous comment $(c)$ and this requirement is crucial in order to proceed. Indeed, Proposition 3.6 transfers and shows that under (4.3) resp. (4.4) the corresponding weighted class $\mathcal{A}^{p}_{(\mathcal{W})}$ resp. $\mathcal{A}^{p}_{\{\mathcal{W}\}}$ is translation invariant. Note (again) that the proofs of Propositions 4.5 and 4.6 require also the non-triviality of the corresponding (smaller) weighted classes and which is not clear by comment $(b)$ above. However, when taking into account the conventions mentioned in Remark 4.4 one is able to conclude in any case.
@@ 6.2. Weighted Fourier image spaces
    Assume that $\{0\}\neq\widehat{\mathcal{A}}^{p}_{[\mathcal{S}]}$ holds and that the underlying space $\mathcal{A}$ is modulation invariant; i.e. if $f\in\mathcal{A}$ then also $g_{x_{0}}:=e^{i\langle\cdot,x_{0}\rangle}f\in\mathcal{A}$ for any $x_{0}\in\mathbb{R}^{d}$. Then the proofs of Propositions 4.5 and 4.6 can be transferred as follows:
[…]
@@ 6.3. Test function spaces appearing in the literature
    Finally, in [23] for $\omega$ satisfying the conditions listed in Definition 1.3 (i.e. PV-weights) the following test function classes have been studied for $U\subseteq\mathbb{R}^{d}$ (see [23, p. 18-19]):
@@
Now focus on the exponential-type system and first note that the equivalences stated in $(c)$ and $(d)$ in Section 5.1 are not valid anymore in this situation: Recall that $(\omega_{1})$ for a non-decreasing radial $\omega$ means $\omega(2t)\leq L\omega(t)+L$ for some $L\geq 1$ and all $t\geq 0$. This constant $L$ yields then a uniform scaling in the parameters of the matrix $\mathcal{W}^{\mathfrak{c}}_{\omega}$; i.e. one infers the relation $L\ell=n$ and which is a problem when restricting the parameters and when $\ell\rightarrow\Lambda^{\prime}$ (resp. $n\rightarrow\Lambda$). Indeed, $L$ yields a transformation (dilation) of the truncation parameter.
@@ 7. On the failure of the convexity condition $(\omega_{4})$
    By (I)(III)(I)-(III), and in view of Remark 2.5 and Lemmas 2.3 and 2.6, we have that $\omega$ is equivalent to a weight function in the sense of Beurling-Björck (see Definition 1.2) and of Petzsche-Vogt (see Definition 1.3).
@@
When combining the information from $(\mathfrak{C}_{1})$ and $(\mathfrak{C}_{2})$ with the characterizations obtained in Theorems 4.1 and 4.2 (see also Proposition 4.9), and by taking into account the comments in Section 5.1, especially the characterization of the growth conditions and relations in comments $(c)$ and $(d)$ there, within the exponential-type weight function matrix setting we obtain: There exist examples of weighted classes treated in the Beurling-Björck-setting which cannot be described (alternatively) by any BMT-weight function and this statement applies to both Roumieu- and Beurling-type spaces.
@@
Finally, note that $\gamma(\omega)=+\infty$ implies $\overline{\gamma}(\omega)=+\infty$ and thus $(\omega_{6})$ fails (recall Section 2.7).
@@ 7.2. Construction of the (counter-)example
Let $\delta=(\delta_{j})_{j\in\mathbb{N}_{>0}}$ be an admissible parameter sequence, arbitrary but from now on fixed. The definition of $\omega$ is depending on chosen $\delta$ which suggests the notation $\varphi_{\omega_{\delta}}$. However, since $\delta$ is now fixed, in order to lighten notation we skip $\delta$ and simply write $\varphi_{\omega}$ except in the very last part Step h: There we verify $(V)$ and hence we have to emphasize the different parameter sequences.
@@
(7.11) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;\varphi_{\omega}(y_{j}):=\frac{2j\varphi_{\omega}(x_{j})}{1-jt_{j}}=\frac{2\delta_{j}j^{2}x_{j}}{1-jt_{j}}=\delta_{j}jy_{j},$
@@
and let us see that this estimate holds: First, $x_{j}\frac{1-jt_{j}}{2}<\frac{x_{j}}{2}$ is clear and so it suffices to have $\left(j-\frac{1}{2}\right)x_{j}\geq j\frac{1-jt_{j}}{2t_{j}}$. Second, by taking into account (7.15) for this estimate it suffices to have $\left(j-\frac{1}{2}\right)\frac{2j}{t_{j}}\geq\frac{j}{2t_{j}}$ and which is clear for all $j\in\mathbb{N}_{>0}$. Thus the claim is verified.
@@
Similarly, for $\ell_{j}$ one has by (7.9), (7.11) and (7.14) that
$\ell_{j}$ $:=\frac{\varphi_{\omega}(x_{j+1})-\varphi_{\omega}(y_{j})}{x_{j+1}-y_{j}}=\frac{\delta_{j+1}(j+1)x_{j+1}-\delta_{j}jy_{j}}{x_{j+1}-y_{j}}=\frac{\delta_{j+1}(j+1)^{2}-\delta_{j}j}{j}.$
@@
and (7.11) implies
$\frac{\varphi_{\omega}(y_{j})}{y_{j}}=\delta_{j}j,$
@@
then the corresponding weights $\omega_{\delta}$ and $\omega_{\delta^{\prime}}$ are not equivalent. From this the assertion follows immediately; see also Remark 7.5 for an explicit one-parameter family.
@@
Let an admissible parameter sequence $\delta$ satisfying $\lim_{j\rightarrow+\infty}\delta_{j}=0$ be given; consider e.g. $\delta_{j}:=\frac{1}{\log(e+j)}$. However, note that this additional condition is not required necessarily in the above arguments and $\delta$ can be bounded by below by some $\epsilon>0$. In any case, under this assumption each power sequence $\delta^{\alpha}=(\delta_{j}^{\alpha})_{j\in\mathbb{N}_{>0}}$, $0<\alpha<1$, is again an admissible parameter sequence since $\delta_{j+1}^{\alpha}\leq\delta_{j}^{\alpha}\leq(j+2)^{\alpha}\delta_{j+1}^{\alpha}\leq(j+2)\delta_{j+1}^{\alpha}$ for all $j\in\mathbb{N}_{>0}$ and $\delta_{j}^{\alpha}\geq\delta_{j}$ for all $j$ such that $\delta_{j}\leq 1$; i.e. for all $j$ sufficiently large. Thus all requirements from (7.5) hold for $\delta^{\alpha}:=(\delta^{\alpha}_{j})_{j\in\mathbb{N}_{>0}}$ and (each) $\delta^{\prime}:=\delta^{\alpha}$ serves as an example for an admissible parameter sequence satisfying the second condition in (7.21); more precisely $\sup_{j\in\mathbb{N}_{>0}}\frac{\delta^{\alpha}_{j}}{\delta^{\beta}_{j}}=+\infty$ for all $0<\alpha<\beta\leq 1$.